In a space 5.82 x 1000000kg rocket experiences a constant force of 6.94 x 1000000000N in the direction of its movement. At time t = 0s,it flies past a planet at a velocity of 57.2m/s. How far away is the rocket from the planet when t = 25.4s?


1
Expert's answer
2022-04-05T19:19:18-0400

The distance, covered by rocket in time t, is


"s=v_0t+\frac{at^2}{2},"


where "v_0" - the initial speed;

a - the acceleration.


From Newton's 2nd law:


"F=ma,\newline
a=\frac{F}{m},"


where F - force, experienced by the rocket.


So,


"s=v_0t+\frac{Ft^2}{2m}=57.2\cdot 25.4+\frac{6.94\cdot 10^9\cdot 25.4^2}{2\cdot 5.82\cdot 10^6}=3.86\cdot10^5\space m."